schoolbusfreak:

A gardener uses 1/3 of a liter of water to water 2/7 of a garden. Watering the entire garden at this rate will require liters of water.

supie:

Still need help?

jhonyy9:

so just use the rule of 3 simple and write in this way 1/3 of a liter of water to water 2/7 of a garden x of a liter of water to water 7/7 of a garden and now cross multiplie to calcule x value x = (7/7 *1/3)/(2/7) = (7/21)/(2/7) = (7/21)*(7/2) = 49/42 = 7/6 = 1 1/6 x = 1 1/6 so what mean that to water the entire garden need more than 1 liter of water hope helped understandably easy